n 4 hours, a man can move 12 blocks, and 3 rocks. In 3 hours, the SAME man can move 4 blocks, and 6 rocks. How long does it take that man to just move 1 block?
Answer by ankor@dixie-net.com(22740)

In 4 hours, a man can move 12 blocks, and 3 rocks.
In 3 hours, the SAME man can move 4 blocks, and 6 rocks.
How long does it take that man to just move 1 block?
:
let b = time (in hrs) to move 1 block
let r = time to move 1 rock
:
Write an equation for each scenario
:
12b + 3r = 4
4b + 6r = 3
:
Multiply the 1st equation by 2, subtract the 2nd equation
24b + 6r = 8
4b + 6r = 3
-------------subtraction eliminates r, find b
20b = 5
b = 5/20
b = 1/4 hr for the man to move one block